### v1

# Extension of the CC(;) Formalism to the Electron Attachment and Ionization Potential Equation-of-Motion Coupled-Cluster Frameworks

Abstract
We combine the electron attachment (EA) and ionization potential (IP) equation-of-motion (EOM) coupled-cluster (CC) approaches with the CC(;) formalism. The resulting methodologies are used to describe the electronic states of several open-shell molecules, with the goal of approximating high-level EA/IP-EOMCC energetics corresponding to a full treatment of 3-particleβ2-hole (3-2) and 3-holeβ2-particle (3-2) excitations on top of CC with singles and doubles (CCSD). We show that the active-orbital-based EA/IP-EOMCC CC(;) approaches, abbreviated as EA/IP-CC(t;3), achieve sub-millihartree accuracies relative to the parent EA-EOMCCSD(3-2)/IP-EOMCCSD(3-2) data using reduced computational effort, while improving upon their completely renormalized EA/IP-CR-EOMCC(2,3) counterparts.
keywords:
Coupled-Cluster Theory , Electron Attachment and Ionization Potential Equation-of-Motion Coupled-Cluster Methods , Completely Renormalized Coupled-Cluster Approaches , CC(;) Formalism1 Introduction
The single-reference coupled-cluster (CC) theory [1, 2] and its equation-of-motion (EOM) [3] extension to electronically excited and electron attached and ionized states are very useful tools for high-accuracy calculations in quantum chemistry. In this Letter, we focus on the electron attachment (EA) and ionization potential (IP) EOMCC approaches [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15], which provide effective and rigorously spin-adapted frameworks for describing electronic states of open-shell species, such as radicals, via electron addition to or electron removal from their closed-shell counterparts, as well as directly accessing molecular ionization and electron attachment spectra. In the EA/IP-EOMCC approaches, the ground () and excited () states of the target ()-electron species are given by
where
is the CC ground state of the underlying -electron system, with representing the cluster operator and designating the -electron reference determinant, which serves as a Fermi vacuum. The electron attaching and ionizing operators, defined as and , respectively, generate the target -electron states via -particleβ-hole [-] or -holeβ-particle [-] excitations out of the -electron ground state . The many-body components of the , , and operators are
and
respectively, where () is the fermionic creation (annihilation) operator associated with spinorbital and we adopt the usual notation in which () denote the spinorbitals occupied (unoccupied) in . The parameters and that specify the truncations in and , respectively, define the conventional hierarchies of EA/IP-EOMCC approximations. Setting and , for example, describes the ()-electron species using up to 2-1 or 2-1 excitations on top of the -electron CC calculations with singles and doubles (CCSD) [16, 17], resulting in the EA-EOMCCSD(2-1) [4, 5, 6] and IP-EOMCCSD(2-1) [9, 7, 8, 6] approaches, respectively. While EA-EOMCCSD(2-1) and IP-EOMCCSD(2-1) are computationally practical approximations, characterized by costs scaling as with the system size , they can only describe open-shell states dominated by 1/1 excitations. They fail to provide meaningful results for states with significant 2-1/2-1 contributions [13, 14, 15, 18, 19, 20, 21]. Higher-level EA/IP-EOMCC approximations, such as EA-EOMCCSD(3-2) [6, 13, 14, 15] and IP-EOMCCSD(3-2) [6, 13, 14, 15], which incorporate the 3-2/3-2 excitations () on top of CCSD (), or their EA-EOMCCSDT(3-2) [6, 10] and IP-EOMCCSDT(3-2) [6, 11] extensions that replace the -electron CCSD state with its more accurate CCSDT [22, 23] counterpart (), are required to accurately describe radical excitation and electron attachment/ioniziation spectra, especially in the presence of significant 2-1/2-1 correlations, which often arise when examining satellite peaks in photoelectron spectra [20, 21] and excited electronic states of radicals [13, 14, 15, 18, 19]. Unfortunately, higher-level EA/IP-EOMCC methods are not easily applicable to larger molecules and basis sets due to the expensive computational costs, which scale as for EA-EOMCCSD(3-2)/IP-EOMCCSD(3-2) and for EA-EOMCCSDT(3-2)/IP-EOMCCSDT(3-2).
The present work aims to address this issue by developing new EA/IP-EOMCC methodologies capable of accurately describing electronic states of radical species at the EA-EOMCCSD(3-2) and IP-EOMCCSD(3-2) levels at small fractions of the computational effort. Several approximate treatments of 3-2/3-2 correlations within the EA/IP-EOMCC frameworks have been explored in the past. Among them are the active-space EA- and IP-EOMCCSDt methods [13, 14, 15], which aim at reducing the costs of EA-EOMCCSD(3-2) and IP-EOMCCSD(3-2) calculations by retaining only the subsets of dominant 3-2 and 3-2 excitations in the and components of selected with the help of active orbitals. Alternatively, one may resort to iterative perturbative treatments [12] or apply noniterative corrections to the low-level EA-EOMCCSD(2-1)/IP-EOMCCSD(2-1) energetics due to 3-2 or 3-2 effects determined with the help of perturbation theory [24, 25, 26]. In this work, we choose yet another strategy and extend the biorthogonal moment expansions defining the CC(;) formalism introduced in Refs.Β [27, 28] (cf.Β Refs.Β [29, 30, 31, 32] for the prior work), which have demonstrated considerable success in the context of particle-conserving CC/EOMCC calculations [27, 28, 33, 34, 35, 36], to the EA/IP-EOMCC regime. This allows us to formulate the noniterative corrections due to all or remaining 3-2 or 3-2 excitations neglected in the EA-EOMCCSD(2-1), IP-EOMCCSD(2-1), EA-EOMCCSDt, and IP-EOMCCSDt calculations without having to rely on traditional perturbation theory concepts. When applied to EA-EOMCCSD(2-1) and IP-EOMCCSD(2-1), the CC(;) corrections developed in this study result in the completely renormalized (CR) EA/IP-EOMCC methodologies abbreviated as EA-CR-EOMCC(2,3) and IP-CR-EOMCC(2,3), respectively, which can be viewed as the EA/IP counterparts of the particle-conserving CR-CC(2,3) [31, 32] and CR-EOMCC(2,3) [32] approaches (cf.Β Ref.Β [37] for the related work in the context of nuclear physics). The proposed CC(;) corrections due to the 3-2 or 3-2 correlations missing in the active-space EA-EOMCCSDt and IP-EOMCCSDt calculations, abbreviated as EA-CC(t;3) and IP-CC(t;3), respectively, are analogous to the CC(t;3) approach of Refs.Β [27, 28]. By examining the ground and excited electronic states of the CH, CNC, C2N, SH, N3, and NCO radicals, we demonstrate that the CC(;)-based EA/IP-CR-EOMCC(2,3) and EA/IP-CC(t;3) corrections significantly improve the corresponding EA-EOMCCSD(2-1)/IP-EOMCCSD(2-1) and EA/IP-EOMCCSDt energetics when compared to the parent EA-EOMCCSD(3-2)/IP-EOMCCSD(3-2) data. In particular, the EA/IP-CC(t;3) approaches developed in this work are capable of recovering the high-level EA-EOMCCSD(3-2)/IP-EOMCCSD(3-2) energetics to sub-millihartree or even microhartree accuracy levels with considerably reduced computational effort.
2 Theory
In developing the EA/IP-CR-EOMCC(2,3) and EA/IP-CC(t;3) approaches, we generalize the CC(;) moment expansions of Refs.Β [27, 28] to the particle-nonconserving EA/IP-EOMCC regime. In the existing particle-conserving CC(;) framework, we solve the CC/EOMCC equations in a subspace of the many-electron Hilbert space called the space and correct the resulting energies for the correlation effects not captured by the -space CC/EOMCC calculations with the help of another subspace of the Hilbert space called the space. In the EA/IP-EOMCC-based formulation of CC(;) proposed in this study, we proceed in a similar manner. Thus, after deciding on the truncation in the cluster operator describing the -electron reference system, we define the space and space as subspaces of the appropriate (-electron sector of the Fock space relevant to the electron attached/ionized wave functions of interest. The space is spanned by the - or - determinants that dominate the target -electron states, whereas the companion space consists of the complementary - or - determinants, which we use to correct the -space EA/IP-EOMCC calculations for missing higher-order correlations.
In the EA/IP-EOMCC methods developed in this work, the underlying -electron reference system is described using CCSD, i.e., , although other truncations in could be considered as well. After solving the CCSD equations for the -electron ground state, we diagonalize the similarity-transformed Hamiltonian in the -electron space, as defined above, to obtain the ground- and excited-state energies of the target electron attached/ionized species and the associated right and left eigenvectors of corresponding to the
and
operators, which define the ket states and their bra counterparts satisfying the biorthonormality condition [the s in Eqs.Β (6) and (7) are elementary - or - excitation operators generating -electron determinants ]. The final CC(;) energies of the -electron species are calculated as
where the noniterative corrections are defined as
with representing the EA/IP-EOMCC analogs of the CC/EOMCC moments considered in Refs.Β [29, 30]. The coefficients multiplying moments in Eq.Β (9) are given by , with defining the corresponding EpsteinβNesbet (EN) energy denominators, although one can also use their MΓΈllerβPlesset (MP) analogs (cf.Β Refs.Β [31, 32, 33]).
This study focuses on the CC(;) approaches aimed at accurately approximating the EA-EOMCCSD(3-2) and IP-EOMCCSD(3-2) energetics at small fractions of the computational costs by correcting the results of the EA/IP-EOMCC calculations with all 1/1, all 2-1/2-1, and, optionally, some 3-2/3-2 excitations on top of CCSD for the many-electron correlation effects associated with all or the remaining 3-2/3-2 contributions using the moment expansions defined by Eq.Β (9). In the case of the noniterative corrections to EA-EOMCCSD(2-1) due to 3-2 correlations, which define the EA-CR-EOMCC(2,3) method, the space is spanned by all 1 () and 2-1 () determinants and the corresponding space consists of all 3-2 determinants (). Similarly, the IP-CR-EOMCC(2,3) approach, which corrects the IP-EOMCCSD(2-1) energetics for 3-2 effects, uses a space spanned by all 1 () and 2-1 () determinants and a space consisting of all 3-2 determinants (). In the case of the active-orbital-based EA-CC(t;3) methodology, the space is defined as all and determinants and the subset of 3-2 determinants selected according to the formula , where the index runs over active unoccupied spinorbitals defined by the user, and the space consists of the remaining 3-2 determinants that are not included in the space. Finally, the IP-CC(t;3) approach adopts a space spanned by all , all , and the subset of 3-2 determinants defined as , where the index denotes the user-defined active occupied spinorbitals, and the space is spanned by the remaining 3-2 determinants.
In analogy to the CC(;) methods of Refs.Β [27, 28, 33, 34, 35, 36] and their CR-CC(2,3) and CR-EOMCC(2,3) predecessors, the EA/IP-CR-EOMCC(2,3) and EA/IP-CC(t;3) approaches offer considerable savings in the computational effort compared to their EA-EOMCCSD(3-2)/IP-EOMCCSD(3-2) parents. This is because the excitation spaces used in the iterative EA/IP-EOMCC steps contain none or only some 3-2/3-2 determinants and the noniterative corrections require fewer CPU operations than a single iteration of EA-EOMCCSD(3-2)/IP-EOMCCSD(3-2). In the next section, we assess how effective the EA/IP-CR-EOMCC(2,3) and EA/IP-CC(t;3) methods are in improving the underlying EA-EOMCCSD(2-1)/IP-EOMCCSD(2-1) and EA/IP-EOMCCSDt computations and recovering the parent EA-EOMCCSD(3-2)/IP-EOMCCSD(3-2) energetics. In doing so, we examine two variants of the EA/IP-CR-EOMCC(2,3) and EA/IP-CC(t;3) approaches, termed EA/IP-CR-EOMCC(2,3)X and EA/IP-CC(t;3)X, respectively, where or . Following the naming convention adopted in Refs.Β [31, 32, 33], the methods labeled with subscript use the EN-like energy denominators in Eq.Β (9) when evaluating corrections . The methods using subscript replace the EN form of by its MP counterpart.
3 Results and Discussion
3.1 Computational Details
To assess the performance of the EA/IP-CR-EOMCC(2,3) and EA/IP-CC(t;3) approaches, we applied them, along with the underlying EA-EOMCCSD(2-1)/IP-EOMCCSD(2-1) and EA/IP-EOMCCSDt methods and their EA-EOMCCSD(3-2)/IP-EOMCCSD(3-2) parents, to the ground and valence excited states of the CH, CNC, C2N, SH, N3, and NCO radical species. The results of our EA-EOMCC calculations for CH, CNC, and C2N are reported in Tables 1 and 2, whereas Tables 3 and 4 summarize the IP-EOMCC computations for SH, N3, and NCO. With the exception of SH, all of our EA/IP-EOMCC calculations correspond to adiabatic transitions; calculations for SH were performed at the ground-state equilibrium geometry. The nuclear geometries characterizing the ground and excited states of CH and the ground-state SβH separation in SH were the same as those used in Ref.Β [15], whereas the CNC, C2N, N3, and NCO molecules were described using the linear geometries optimized with the EA (CNC and C2N) or IP (N3 and NCO) variants of the symmetry-adapted-cluster (SAC) configuration interaction (CI) approach [38] abbreviated as SDT- [39] in Ref.Β [18]. All EA-EOMCC calculations for CH, CNC, and C2N reported in this work employed the restricted HartreeβFock (RHF) orbitals of the underlying (CH)+, (CNC)+, and (C2N)+ cationic references, whereas the IP-EOMCC calculations for SH, N3, and NCO used the RHF orbitals of the closed-shell (SH)-, (N3)-, (NCO)- anions. Our calculations for CH employed the aug-cc-pVQZ basis set [40, 41], while for SH, we used the aug-cc-pV(T+d)Z [42] basis for sulfur and aug-cc-pVTZ [40, 41] for hydrogen. Following Ref.Β [18], the CNC, C2N, N3, and NCO radicals were described using the DZP[4s2p1d] basis set [43, 44]. The active orbital spaces used to define the and components of entering the EA/IP-EOMCCSDt and subsequent EA/IP-CC(t;3) calculations were the same as those used in Refs.Β [15, 18]. Thus, the EA-EOMCCSDt and EA-CC(t;3) calculations for CH used an active space consisting of the lowest-energy orbitals unoccupied in (CH)+, whereas the active spaces employed in the analogous calculations for CNC and C2N were spanned by the four orbitals corresponding to the two lowest shells unoccupied in (CNC)+ and (C2N)+. In the case of the IP-EOMCCSDt and IP-CC(t;3) computations for SH, N3 and NCO, the active spaces consisted of the highest-energy shell occupied in the respective parent (SH)-, (N3)-, and (NCO)- anions. The lowest-energy orbitals corresponding to the 1s shells of carbon, oxygen, nitrogen, and sulfur atoms were frozen in all post-RHF steps.
All of the EA- and IP-EOMCC methods used in our calculations, including the new EA-CR-EOMCC(2,3), IP-CR-EOMCC(2,3), EA-CC(t;3), and IP-CC(t;3) approaches developed in this work and the EA-EOMCCSD(2-1), IP-EOMCCSD(2-1), EA-EOMCCSDt, IP-EOMCCSDt, EA-EOMCCSD(3-2), and IP-EOMCCSD(3-2) methods, were implemented in the open-source CCpy package available on GitHub [45]. The required RHF orbitals and one- and two-electron molecular integrals were generated with GAMESS [46]. GAMESS, especially its EA-EOMCCSD(2-1), IP-EOMCCSD(2-1), EA-EOMCCSDt, IP-EOMCCSDt, EA-EOMCCSD(3-2), and IP-EOMCCSD(3-2) routines developed in Ref.Β [15], was also useful in testing our CCpy codes. Our EA- and IP-CR-EOMCC(2,3) routines in CCpy were also tested against their automated implementations obtained with pβ q [47].
Our selection of the above six open-shell species is motivated by the previous studies reported in Refs.Β [15, 18], which demonstrated that EA-EOMCCSD(3-2) and IP-EOMCCSD(3-2) β the parent approaches for all of the CC(;) schemes considered in this study β can provide electronic spectra that closely match the results obtained with other high-level methods, including multireference CI (CH and SH) and EA/IP SAC-CI SDT- (CNC, C2N, N3, and NCO). With the exception of the state of C2N, which is characterized by significant 3-2 correlations [18], the EA-EOMCCSD(3-2)/IP-EOMCCSD(3-2) calculations also predict adiabatic excitation energies for the low-lying states of CH, CNC, C2N, N3, and NCO in good agreement with experiment, with errors ranging from 0.03 to 0.50 eV and an average deviation of 0.20 eV (cf.Β Refs.Β [15, 18]). These observations justify EA-EOMCCSD(3-2) and IP-EOMCCSD(3-2) as benchmarks for evaluating the accuracy of their CC(;)-based approximations.
3.2 Performance of the CC(;)-based EA/IP-EOMCC Methodologies
We begin our discussion by examining the EA-EOMCC calculations reported in Tables 1 and 2. It is clear that while the basic EA-EOMCCSD(2-1) approach can provide qualitative results for the ground states of CH, CNC, and C2N, which are characterized by errors relative to EA-EOMCCSD(3-2) of 2.173, 4.881, and 4.696 millihartree, respectively, it fails to describe the valence excited states, producing errors on the order of 0.1 hartree. The inability of EA-EOMCCSD(2-1) to balance the description of ground and excited states of CH, CNC, and C2N results in highly inaccurate adiabatic excitation spectra, which deviate from their EA-EOMCCSD(3-2) counterparts by 2β4 eV [15, 18]. The failure of EA-EOMCCSD(2-1) is largely a consequence of neglecting the 3-2 effects, which are essential for accurately describing electronic states with significant 2-1 character. This can be mitigated by the noniterative EA-CR-EOMCC(2,3) corrections. As shown in Tables 1 and 2, the 2β5 millihartree differences between the EA-EOMCCSD(2-1) and EA-EOMCCSD(3-2) energies of the ground states of CH, CNC, and C2N are reduced by a factor of 2β3 when the EA-CR-EOMCC(2,3) corrections to EA-EOMCCSD(2-1) are employed. In the case of excited states, the improvements offered by the EA-CR-EOMCC(2,3) corrections are even more substantial. The massive errors relative to EA-EOMCCSD(3-2) obtained with EA-EOMCCSD(2-1), which are 69.816β117.876 millihartree for CH, 112.241β118.827 millihartree for CNC, and 111.733β158.282 millihartree for C2N, reduce in the EA-CR-EOMCC(2,3)A calculations to 17.332β31.850, 20.996β22.060, and 20.003β29.695 millihartree, respectively. The analogous errors characterizing the EA-CR-EOMCC(2,3)D calculations, which are 0.115β12.505 millihartree for CH, 14.346β14.517 millihartree for CNC, and 18.601β33.535 millihartree for C2N, are similar or smaller. The noniterative EA-CR-EOMCC(2,3)A and EA-CR-EOMCC(2,3)D corrections also improve the adiabatic excitation spectra resulting from the EA-EOMCCSD(2-1) calculations, reducing the 2β4 eV errors relative to EA-EOMCCSD(3-2) obtained with EA-EOMCCSD(2-1) to 0.45β0.84 and 0.03β0.97 eV, respectively. While the EA-CR-EOMCC(2,3) methodology helps temper the failures observed in the EA-EOMCCSD(2-1) calculations, the 20β30 millihartree errors relative to EA-EOMCCSD(3-2) characterizing the majority of our EA-CR-EOMCC(2,3) results remain. In analogy to the particle-conserving CC(;) framework [27, 28, 33, 34, 35, 36], this can be attributed to the decoupling of the component of from its lower-rank and counterparts when the EA-CR-EOMCC(2,3) corrections are constructed. As shown in Tables 1 and 2, this problem can be addressed by turning to the EA-CC(t;3) approach, which replaces EA-EOMCCSD(2-1) by EA-EOMCCSDt prior to determining the 3-2 corrections.
Upon examining the EA-EOMCCSDt and EA-CC(t;3) data reported in Tables 1 and 2, the benefits offered by including the dominant 3-2 excitations selected via active orbitals in the iterative EA-EOMCC steps are readily apparent. The energies resulting from the EA-EOMCCSDt calculations for the ground and excited states of CH, CNC, and C2N differ from their EA-EOMCCSD(3-2) parents by 1.454β3.243, 1.414β1.910, and 0.543β1.606 millihartree, respectively, which in most cases is a massive improvement over EA-EOMCCSD(2-1). In fact, with the exception of the and states of CH and the state of CNC, EA-EOMCCSDt outperforms EA-CR-EOMCC(2,3) as well, which highlights the importance of incorporating the coupling between and its lower-rank and counterparts when treating states characterized by stronger 2-1 correlations. The adiabatic excitation spectra of CH, CNC, and C2N computed with EA-EOMCCSDt are also very accurate, with errors relative to full EA-EOMCCSD(3-2) not exceeding 0.05 eV (cf.Β Refs.Β [15, 18]). By correcting the EA-EOMCCSDt energetics for the remaining 3-2 correlations using EA-CC(t;3), we obtain a virtually perfect agreement with the parent EA-EOMCCSD(3-2) data. Indeed, when using the EA-CC(t;3)A approach, the 1β3 millihartree errors relative to EA-EOMCCSD(3-2) resulting from the EA-EOMCCSDt calculations for CH, CNC, and C2N reduce to 0.221β0.412, 0.194β0.241, and 0.084β0.201 millihartree, respectively. The EA-CC(t;3)D corrections to EA-EOMCCSDt, which use the EN denominators in calculating the coefficients in Eq.Β (9) instead of their MP counterparts, are even more accurate, reproducing the target EA-EOMCCSD(3-2) energetics to within tens of microhartree or less. Even the largest deviation of 0.106 millihartree, observed in the EA-CC(t;3)D calculations for the state of CH, is still an order of magnitude smaller than the error obtained with EA-EOMCCSDt. It is important to emphasize that the EA-CC(t;3) calculations reported in Tables 1 and 2 are much less expensive than the parent EA-EOMCCSD(3-2) computations, while providing virtually identical energetics. For example, the spaces adopted in our EA-EOMCCSDt and EA-CC(t;3) computations for CNC and C2N used only about 30% of all 3-2 determinants. In the case of CH, the fractions of 3-2 determinants included in the corresponding spaces were even smaller (less than 5% of all 3-2 determinants). As demonstrated in Ref.Β [15], the use of small fractions of 3-2 determinants in the iterative EA-EOMCCSDt steps translates into substantial savings in the computational costs relative to the parent EA-EOMCCSD(3-2) approach. The EA-CC(t;3) corrections to EA-EOMCCSDt do not change this, since they involve fewer CPU operations than a single iteration of EA-EOMCCSD(3-2).
Similar accuracy patterns are observed in the results of the CC(;)-based IP-EOMCC calculations reported in Tables 3 and 4. Much like EA-EOMCCSD(2-1), the basic IP-EOMCCSD(2-1) approach is generally incapable of providing an accurate description, especially when 2-1 effects become more substantial [15, 18]. In the case of SH, the discrepancies between the IP-EOMCCSD(2-1) and parent IP-EOMCCSD(3-2) energetics are on the order of 0.1β0.2 hartree for all but the and states, which are described more accurately by IP-EOMCCSD(2-1) since these two states are the only ones in Table 3 dominated by 1 correlations. The same is true when examining the vertical excitation spectrum of SH obtained with IP-EOMCCSD(2-1), where the errors relative to IP-EOMCCSD(3-2) are 0.06 eV for the state and 3.5β5.3 eV for the remaining excited states shown in Table 3 (cf.Β Ref.Β [15]). In the case of the electronic states of N3 and NCO reported in Table 4, for which 2-1 effects are small, the IP-EOMCCSD(2-1) method fares better, providing adiabatic transition energies in good agreement (0.3 eV or less) with the IP-EOMCCSD(3-2) results [18], but this is partly due to a fortuitous cancellation of the 10β20 millihartree errors characterizing the IP-EOMCCSD(2-1) energies relative to IP-EOMCCSD(3-2). In analogy to the EA-EOMCC computations, we can improve the IP-EOMCCSD(2-1) results by applying the IP-CR-EOMCC(2,3) corrections due to 3-2 effects. Indeed, as shown in Table 3, the 9.898 and 12.149 millihartree errors relative to IP-EOMCCSD(3-2) obtained with IP-EOMCCSD(2-1) for the and states of SH reduce in the IP-CR-EOMCC(2,3)A calculations to 2.904 and 4.449 millihartree, respectively. The IP-CR-EOMCC(2,3)D approach reduces these errors even further, to 1.576 and 3.201 millihartree. The error reductions offered by the IP-CR-EOMCC(2,3) corrections are even larger when the remaining excited states of SH considered in Table 3, characterized by significant 2-1 contributions, are examined. For these states, the errors relative to IP-EOMCCSD(3-2) obtained with IP-EOMCCSD(2-1), which range from 137.280 to 203.145 millihartree, reduce in the IP-CR-EOMCC(2,3)A and IP-CR-EOMCC(2,3)D calculations to 0.359β39.214 and 3.201β64.470 millihartree, respectively. As a result, the vertical excitation energies corresponding to the state of SH obtained with the IP-CR-EOMCC(2,3) corrections agree with their IP-EOMCCSD(3-2) parent value to within 0.04 eV and the 3.5β5.3 eV differences between the IP-EOMCCSD(2-1) and IP-EOMCCSD(3-2) excitation energies characterizing the remaining excited states of SH shown in Table 3 decrease in the IP-CR-EOMCC(2,3)A and IP-CR-EOMCC(2,3)D computations to 0.07β0.99 eV and 0.14β1.71 eV, respectively. For N3 and NCO, the improvements offered by the 3-2 corrections to IP-EOMCCSD(2-1) are less dramatic, since the electronic states of these two molecules included in Table 4 are dominated by 1 excitations out of (N3)- and (NCO)-, but the reductions in errors relative to IP-EOMCCSD(3-2), from 9.587β20.154 millihartree obtained with IP-EOMCCSD(2-1) to 0.808β7.044 millihartree obtained with variants A and D of IP-CR-EOMCC(2,3), are still quite substantial. We can conclude that the IP-CR-EOMCC(2,3) corrections are helpful, but, as shown in Tables 3 and 4, they are not as effective in approximating the IP-EOMCCSD(3-2) energetics as the IP-CC(t;3) corrections to IP-EOMCCSDt.
The IP-CC(t;3) results reported in Tables 3 and 4 and their IP-EOMCCSDt counterparts clearly show the benefits of including the leading 3-2 determinants in the relevant spaces prior to determining the 3-2 corrections. Indeed, with the exception of the state of NCO, the IP-EOMCCSDt approach substantially improves the IP-EOMCCSD(2-1) energetics, reducing the 9.898β203.145, 13.078β14.623, and 9.587β20.154 millihartree errors relative to IP-EOMCCSD(3-2) obtained for SH, N3, and NCO with IP-EOMCCSD(2-1) to 0.015β1.003, 1.400β6.215, and 1.926β2.205 millihartree, respectively. In the case of the state of NCO, the improvement over the basic IP-EOMCCSD(2-1) level offered by IP-EOMCCSDt is minimal, but this is no longer an issue when the IP-CC(t;3)A and IP-CC(t;3)D corrections to IP-EOMCCSDt are employed. The results in Tables 3 and 4 clearly show that the IP-CC(t;3) corrections not only improve the IP-EOMCCSDt energetics, but they are also considerably more accurate than their IP-EOMCCSD(2-1)-based IP-CR-EOMCC(2,3) counterparts, and this applies to all of the electronic states of SH, N3, and NCO considered in this work. For SH, both IP-CC(t;3)A and IP-CC(t;3)D recover the parent IP-EOMCCSD(3-2) data to within small fractions of a millihartree. In the cases of N3 and NCO, the IP-CC(t;3)A method shows a maximum error relative to IP-EOMCCSD(3-2) of just 1.690 millihartree, while the more complete IP-CC(t;3)D approach recovers its parent energetics to within one millihartree for all electronic states. In analogy to the EA-CC(t;3) approach, the IP-CC(t;3) calculations provide IP-EOMCCSD(3-2)-level results with reduced computational costs. Indeed, the spaces entering our IP-EOMCCSDt and IP-CC(t;3) calculations for SH, N3, and NCO include approximately 60% of all 3-2 determinants, which translate to sizeable speedups compared to the full IP-EOMCCSD(3-2) computations. In fact, the iterative IP-EOMCCSDt and noniterative IP-CC(t;3) steps require CPU efforts that are comparable to the underlying -electron CCSD calculations.
4 Summary
In this work, we have developed the CC(;)-based EA/IP-EOMCC approaches that correct the energetics resulting from the EA-EOMCCSD(2-1) and IP-EOMCCSD(2-1) calculations as well as their active-space EA/IP-EOMCCSDt extensions for all or the remaining 3-2 or 3-2 correlation effects. While the EA/IP-CR-EOMCC(2,3) methods provide sizeable improvements over the low-level EA-EOMCCSD(2-1) and IP-EOMCCSD(2-1) data, the most accurate results are obtained using active-orbital-based EA/IP-CC(t;3) calculations, which are capable of recovering the parent EA-EOMCCSD(3-2)/IP-EOMCCSD(3-2) energetics to within a fraction of a millihartree using reduced computational effort. In the future, it would be interesting to extend the present CC(;) formalism to target higher-level EA/IP-EOMCC approaches, such as EA-EOMCCSDT(3-2) and IP-EOMCCSDT(3-2), in a computationally efficient manner. This may be accomplished, for example, by replacing the -electron CCSD similarity-transformed Hamiltonian with its more accurate CCSDt [48, 49, 50] counterpart. Furthermore, we can combine the EA/IP-EOMCC-based CC(;) framework with the semi-stochastic [33], selected-CI-based [35], and adaptive CC(;) [36] techniques, to automate the construction of the and spaces entering the CC(;) computations.
CRediT authorship contribution statement
Karthik Gururangan: Methodology, Software, Data curation, Formal analysis, Validation, Writing - original draft. Stephen H. Yuwono: Methodology, Software, Data curation, Formal analysis, Validation, Writing - original draft. A. Eugene DePrince III: Funding acquisition, Project administration, Resources, Supervision, Writing - reviewing and editing. Piotr Piecuch: Conceptualization, Methodology, Formal analysis, Investigation, Funding acquisition, Project administration, Resources, Supervision, Validation, Writing - reviewing and editing.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
The data that support the findings of this study are available within the article.
Acknowledgments
This work has been supported by the Chemical Sciences, Geosciences and Biosciences Division, Office of Basic Energy Sciences, Office of Science, U.S. Department of Energy (Grant No. DE-FG02-01ER15228 to P.P). A.E.D. acknowledges support by the Office of Advanced Scientific Computing Research and Office of Basic Energy Sciences, Scientific Discovery through the Advanced Computing (SciDAC) program, Office of Science, U.S. Department of Energy, under Award No. DE-SC0022263.